# A common first integral from three-body secular theory and Kepler billiards

### v1

@@ Abstract.
We observe that a particular first integral of the partially-averaged system in the secular theory of the three-body problem appears also as an important conserved quantity of integrable Kepler billiards. In this note we illustrate their common roots with the projective dynamics of the two-center problem. We then combine these two aspects to define a class of integrable billiard systems on surfaces of constant curvature.
@@ 1. Introduction
In a Euclidean space of dimension at least $2$, the Kepler problem is super-integrable: It possesses more independent first integrals for the system to be integrable in the sense of Liouville. Concretely, in addition to the total energy and the angular momentum, additional conserved quantities are provided by the components of the Laplace-Runge-Lenz vector. Dynamically, this implies that the system is properly-degenerate: All bounded, non-singular orbits are closed. Out of $d$-frequencies (in a proper set of Action-Angle coordinates), only $1$-frequency is non-zero. This proper-degeneracy of the Kepler problem is the main reason for many degeneracies in celestial mechanics. This makes it hard to study perturbations, but also leaves the possibility of considering systems with varying Kepler orbits, in a discrete or a continuous way, without destroying the integrability. Moreover, this can even be made with fixed Keplerian energy.
@@
serves as a conserved quantity for two types of these systems recently studied, in which $C$ is the total angular momentum, $A_{1}$ is a component of the Laplace-Runge-Lenz vector, and $h$ is a parameter:
  • Keplerian orbits evolving under a partially-averaged Newtonian potential, in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. This has been noticed by the first author in [16]. In this situation, up to certain normalization, $h$ denotes half of the distance between the origin and the position of the second, motionless particle.
@@
The existence of $D$ in the study of Gallavotti-Jauslin [9] is used to disprove an ergodic assertion in a paper of Boltzmann [2]. This has intrigued many studies ever since [10, 25]. The function $D$ is shown to hold as well for Kepler billiard systems with a branch of conic section as reflection wall, provided that the Kepler center lies at a focus of the conic section, $h$ is the focus-center distance of the conic section, and $A_{1}$ the component of the Laplace-Runge-Lenz vector along the axis of foci [22, 23].
@@
A main observation of this article is that the appearance of $D$ in both of systems can be understood in terms of projective dynamics [1]. Consequently we make several further generalizations:
@@
2. The two-center problems
2.1. Two center problems on constant curvature spaces
We consider the two-center problem in an Euclidean space $\mathbb{R}^{d}$. In this system, a moving particle of unit mass moves in the space $\mathbb{R}^{d}$ under the attractions of two fixed Kepler centers. This system is an integrable system, already known to Euler [4] and Lagrange [11] for $d=2,3$. Indeed rotations around the axis of the centers provides $d-2$ first integrals commuting with the energy in $\mathbb{R}^{d}$ by Noether’s theorem, and thus the integrability of the system follows from one additional non-trivial first integral, which has been identified by these authors.
This type of problems can be imposed on the sphere as well. The spherical 2-center problem describes the motion of a particle moving on $\mathbb{S}^{d}$ attracted by two pairs of antipodal centers of Kepler-Coulomb type, generated from potentials proportional to the cotangent of the central angle. The identification of such 1-center potentials of Kepler-Coulomb type is made to preserve the Kepler law on a spherical universe as much as possible, in particular to preserve the elliptic law. This was done by Graves [8] and Serret [21]. The two-center problem in the hyperbolic 3-space has been studied by Killing [13].
@@ 2.2. Projective Dynamics of the Two-Center/Lagrange Problems
In [1], Albouy gave an explanation that the additional non-trivial first integral of the two center problem in $\mathbb{R}^{d}$ is ultimately related to the energy of a corresponding system defined on the sphere, which is the spherical two center problem on a hemisphere of $\mathbb{S}^{d}$.
We recall the construction more precisely. We consider $\mathbb{S}^{d}$ embedded as unit sphere in $\mathbb{R}^{d+1}$, with $\mathbb{R}^{d}$ embedded as the hyperplane $\mathbb{R}^{d}\times\{-1\}\subset\mathbb{R}^{d+1}$. The lower hemisphere $\mathbb{S}^{d}_{SH}:=\{(x_{1},\cdots,x_{d+1})\in\mathbb{S}^{d},x_{d+1}<0\}$ of $\mathbb{S}^{d}$ is thus projected to $\mathbb{R}^{d}$ via central projection from the origin in $\mathbb{R}^{d+1}$. If the centers of the two center problem are put symmetrically with respect to the contact point of $\mathbb{S}^{d}$ with $\mathbb{R}^{d}\times\{-1\}$, then this projection also project unparametrized orbits of this spherical problem to unparametrized orbits of the two-center problem in $\mathbb{R}^{d}$, defined with respect to a modified metric $\|\cdot\|_{a}$ given below [1]. The spherical energy thus induces an additional first integral for the two-center problem in $\mathbb{R}^{d}$ and vice versa, the energy of the Euclidean two-center problem is an additional first integral for the spherical two-center problem. Moreover, we may as well put a harmonic potential in the middle of the two Kepler centers and all these hold for the modified system, called the Lagrange problem [11] as well. This can as well be defined on $\mathbb{S}^{d}_{SH}$, with a potential proportional to the squared tangent of the central angle as the spherical harmonic potential.
We denote the masses of the two Kepler centers and the string constant of the Hooke center in $\mathbb{R}^{d}$ as $m_{1},m_{2},f$ respectively. We put the centers in $\mathbb{R}^{d}$ at $(0,\cdots,0,\pm a)$. In $\mathbb{R}^{d}$, the systems are defined with respect to the affine norm $\|(v_{1},\cdots,v_{d})\|_{a}:=\sqrt{v_{1}^{2}+\cdots+v_{d-1}^{2}+\frac{v_{d}^{2}}{1+a^{2}}}$.
@@
These systems, as well as those defined on $\mathbb{S}^{d}_{SH}$ are thus integrable. Similar result holds between systems defined in $\mathbb{S}^{d}$ and in $\mathbb{H}^{d}$ as well. We refer to [22] for precise statements.
For the problem in $\mathbb{R}^{d}$, we denote the energy of the problem as $E_{Eucl}$. We consider $\mathbb{R}^{d}$ as a chart for the lower hemi-sphere $\mathbb{S}_{SH}$ and we denote by $E_{sph}$ the spherical energy in terms of positions and velocities, so as a function on the tangent bundle $T\mathbb{R}^{d}$. Note that Legendre transformation allows to identify the tangent and cotangent bundles. In the sequel we shall also regard them as functions on cotangent bundle when necessary.
@@ 2.3. Relationship among first integrals for the two center problem
We consider the two center problem and argue in the case $d=2$. Due to the rotation invariance the formulas hold for any $d\geq 2$ as well.
We put the Kepler centers at $(0,a)$ and $(0,-a)$ in $\mathbb{R}^{2}$, with masses $m_{1}$ and $m_{2}$ respectively. We equip $\mathbb{R}^{2}$ with the norm $\|\cdot\|_{a}$, defined as $\|(u,v)\|_{a}=\sqrt{u^{2}+\frac{v^{2}}{1+a^{2}}}$.
This is the same setting as in [25]. Same argument therein, by just adding the potential of the second center, leads to
@@
We use the affine coordinate change $(\xi,\eta)\mapsto(x=\xi,y=\sqrt{1+a^{2}}\eta+a)$ to bring the norm on $\mathbb{R}^{2}$ into the standard Euclidean one. With this change we have put the second center at $(0,-2a/\sqrt{1+a^{2}})$ in this new coordinates. And we have
@@
Further set $E_{Kep}:=E_{Eucl}|_{m_{2}=0}$ the Keplerian energy with respect to the first center. Then
@@
in which $D$ as appeared in (1) is a first integral of $E_{Kep}$. The term $V_{Eucl,2}$ is linear in $m_{2}$, as well as $K$:
@@
We have
(5) $\{E_{Eucl},E_{sph}\}=0,$
since $E_{sph}$ is invariant under the flow of $E_{Eucl}$.
@@ 2.4. Averaging in the two center problem
We switch to the cotangent bundle while keeping the notations for the functions unchanged. We argue on the region on which $E_{Kep}$ defines a circle action (i.e. region on which all the orbits are closed) for which the orbits in $\mathbb{R}^{d}$ avoid the second Kepler center. Denote the set of orbits of $E_{Kep}$ in $\mathbb{T}^{*}\mathbb{R}^{d}$ whose projection passes through the second Kepler center by $\mathcal{I}$. We shall exclude this set as $V_{Eucl,2}$ encounters singularities along these orbits. Then the region $\mathbb{R}$ defined by having negative Keplerian energy $E_{Kep}<0$ and non-zero Kepler angular momentum $C_{Kep}\neq 0$ to exclude collisional orbits in $T^{*}\mathbb{R}^{d}\setminus\mathcal{I}$:
:=$\mathcal{R}:=\{E_{Kep}<0,C_{Kep}\neq 0\}\subset T^{*}\mathbb{R}^{d}\setminus\mathcal{I}.$
Moreover, the latter condition excluding collision orbits can likely be dropped since averaging is well-defined up to collisional orbits by mean of regularization. In dimension 2 and 3 this can be justified using methods and coordinates explained in [6] and [27] respectively. This is likely the case for all dimensions but we shall not deal with this issue here.
@@
[…]
in which the integration is understood as averaging over closed orbits of $E_{Kep}$.
The definition does not require the choice of a set of canonical coordinates. When $d=2,3$, several sets of coordinates are available, including the classical Delaunay variables and Poincaré coordinates. Moser constructed action-angle coordinates of Delaunay type for $\mathbb{R}^{d}$ [15]. In general it is not always easy to construct these coordinates, which is even more complicated in the case when more bodies were involved [19]. Moreover, all these coordinates become singular at certain types of orbits.
Nevertheless, in a trivialization of this circle bundle, we can have an (abstract) set of canonical coordinates $(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$ after a further localization on the space of closed Keplerian orbits, in which $L$ is the circular angular momentum as in the Delaunay coordinates, and $\tilde{\ell}$ is an angle parametrizing the Keplerian orbits in this trivialization (shifted from the mean anomaly by a phase, when both are well-defined), and $(u_{i},v_{i})$ are secular variables, in the sense that they are functions of the orbits only, independent of the actual position and velocity of the particle on the orbit. We refer to [26] for a detailed discussion on these abstract local coordinates. It is not hard to see that Definition 2.1 is independent on the choice of local coordinates. Thus averaging is well-defined in a non-trivial principal $S^{1}$-bundle, in our case defined by motions along closed Keplerian orbits.
The following proposition is a coordinate-free formulation of a main observation in [16]. Without specifying the coordinates used in the announcement we bring a small generalization.
@@
We cover $M$ by open submanifolds on which the above-mentioned canonical coordinates are defined. Let $\tilde{M}$ be any of these submanifolds, with local canonical coordinates $(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$.
@@
Then
$E_{Eucl}=E_{Eucl,a}+V_{pl,2}-\langle V_{pl,2}\rangle_{\tilde{\ell}}.$
We apply averaging theory. There is only one frequency that we need to average. Moreover we may use $m_{2}$ to play the role of a small parameter. In this setting, we know that there is a canonical transformation $\phi:\tilde{M}\mapsto\tilde{M}$ of order $O(m_{2})$ which conjugate $E_{Eucl}$ to $E_{Eucl,a}$, that is to say,
$E_{Eucl,a}=E_{Eucl}\circ\phi.$
And
$\{E_{Eucl,a},D\circ\phi+K\circ\phi\}=0.$
@@
$\{E_{Eucl,a},D\circ\phi+K\circ\phi\}=0=\{E_{Eucl,a},D\}+O(m_{2}).$
Since the above equation holds for all $m_{2}$, we have
$\{E_{Eucl,a},D\}=0.$

### v2

@@ Abstract.
A particular first integral of the partially-averaged system in the secular theory of the three-body problem appears also as an important conserved quantity of integrable Kepler billiards. In this note we illustrate their common roots with the projective dynamics of the two-center problem. We then combine these two aspects to define a class of integrable billiard systems on surfaces of constant curvature.
@@ 1. Introduction
In a Euclidean space of dimension at least $2$, the Kepler problem is super-integrable: It possesses more independent first integrals for the system to be integrable in the sense of Liouville. Concretely, in addition to the total energy and the angular momentum, additional conserved quantities are provided by the components of the Laplace-Runge-Lenz vector. Dynamically, this implies that the system is properly-degenerate: All bounded, non-singular orbits are closed. Out of $d$-frequencies (say, in a proper set of Action-Angle coordinates), only $1$-frequency is non-zero. This proper-degeneracy of the Kepler problem is the main reason for many degeneracies in celestial mechanics. The perturbation theory to Kepler problems is therefore hard to investigate. Nevertheless, this also leaves the possibility of considering systems with varying Kepler orbits, in a discrete or a continuous way, without destroying the integrability, even with fixed Keplerian energy.
@@
serves as a conserved quantity in addition to the energy for two types of these systems recently studied, in which $C$ is the total angular momentum, $A_{1}$ is a component of the Laplace-Runge-Lenz vector, and $h$ is a parameter:
  • Keplerian orbits evolving under a partially-averaged Newtonian potential, in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. This has been noticed by the first author in [19]. In this situation, up to certain normalization, $h$ denotes half of the distance between the origin and the position of the second, motionless particle. $A_{1}$ is taken as the component of the Laplace-Runge-Lenz vector in the direction of the second particle.
@@
The existence of $D$ in the study of Gallavotti-Jauslin [11] is used to disprove an ergodic assertion in a paper of Boltzmann [3]. This has intrigued many studies ever since [7, 12, 29]. The function $D$ is shown to hold as well for Kepler billiard systems with a branch of conic section as reflection wall, provided that the Kepler center lies at a focus of the conic section, $h$ is the focus-center distance of the conic section, and $A_{1}$ the component of the Laplace-Runge-Lenz vector along the axis of foci [26, 27].
@@
A main observation of this article is that the appearance of $D$ in both of systems can be understood in terms of projective dynamics [1]. See also [23] in which we present an alternative computation with full details. We make several further generalizations:
@@
2. The Kepler and two-center problems
2.1. Kepler and two center problems on constant curvature spaces
We consider the two-center problem in an Euclidean space $\mathbb{R}^{d}$. In this system, a moving particle of unit mass $q$ moves in the space $\mathbb{R}^{d}$ under the Newtonian attractions of two fixed Kepler centers $Z_{1},Z_{2}\in\mathbb{R}^{d}$. If the masses of the centers are $m_{1}$ and $m_{2}$ respectively, then the equations of motion is
$\ddot{q}=-\nabla_{q}V(q),$
with the potential
$V(q)=-\dfrac{m_{1}}{\|Z_{1}-q\|}-\dfrac{m_{2}}{\|Z_{2}-q\|}.$
This system is an integrable system, already known to Euler [6] and Lagrange [13] for $d=2,3$. Indeed by Noether’s theorem, rotations around the axis of the centers provide d2d-2 first integrals commuting with the energy in $\mathbb{R}^{d}$. The integrability of the system follows from one additional non-trivial first integral, which has been identified by these authors.
[…]
$\nabla_{\dot{q}}\dot{q}=-\hbox{grad}_{g_{0}}V,$
in which both the covariant derivative and the gradient is defined with respect to the round metric $g_{0}$ on $\mathbb{S}^{d}$. This is a natural mechanical system, and the total energy
$\dfrac{1}{2}g_{q}(\dot{q},\dot{q})+V(q)$
is a conserved quantity.
As noted by Graves [10] and Serret [25], the natural extension of Kepler problem from the Euclidean space to the (unit) sphere which preserves the elliptic law is given by a potential proportional to the cotangent of the central angle the moving particle made with the Kepler center. The spherical Kepler potential respect to a center $Z_{1}\in\mathbb{S}^{d}$ of mass $m_{1}$ is thus given by $m_{1}\cot{\theta_{1}}$. The angle $\theta_{1}$ is the central angle the particle $q$ made from $Z_{1}$. Note that the antipodal point of $Z_{1}$ on $\mathbb{S}^{d}$ becomes another singularity for the potential, repulsive if $Z_{1}$ is attractive and attractive if $Z_{1}$ is repulsive, so Kepler singularities on the sphere appear in pairs. It is known to these authors that all non-singular orbits are spherical ellipses, which presents the form of elliptic law on the sphere. Further study shows that there are even more similarities of the Euclidean and spherical Kepler problems [17, 2].
As in the Euclidean case, the spherical 2-center problem describes the motion of a particle moving on $\mathbb{S}^{d}$ attracted by two pairs of antipodal centers of Kepler type. On the sphere $S^{d}$ this system is thus described by a potential of the form $m_{1}\cot{\theta_{1}}-m_{2}\cot{\theta_{2}}$. This is an integrable system closely related to the Euclidean two-center problem [1], and will be important for this note.
We may define Kepler problem and the two-center problems in the hyperbolic spaces in an analogous manner. They share many similarities with their spherical and Euclidean analogues. We note that the two-center problem in the hyperbolic 3-space has been first studied by Killing [15].
@@ 2.2. Projective Dynamics of the Two-Center/Lagrange Problems
In [1], Albouy gave an explanation that the additional non-trivial first integral of the two center problem in $\mathbb{R}^{d}$ is ultimately related to the energy of a corresponding system defined on the sphere, which is the two-center problem on a hemisphere of $\mathbb{S}^{d}$. It is shown that the natural central projection from the center of $\mathbb{S}^{d}$ projects unparametrized orbits of one system to unparametrized orbits of the other and thus both energies of these two systems give rise to a pair of independent, commuting first integrals for both systems.
Here we recall the construction. We consider $\mathbb{S}^{d}$ embedded as unit sphere in $\mathbb{R}^{d+1}$, with $\mathbb{R}^{d}$ embedded as the hyperplane $\mathbb{R}^{d}\times\{-1\}\subset\mathbb{R}^{d+1}$. The lower hemisphere $\mathbb{S}^{d}_{SH}:=\{(x_{1},\cdots,x_{d+1})\in\mathbb{S}^{d},x_{d+1}<0\}$ of $\mathbb{S}^{d}$ is thus projected to $\mathbb{R}^{d}$ via central projection from the origin in $\mathbb{R}^{d+1}$. If the centers of the two center problem are put symmetrically with respect to the contact point of $\mathbb{S}^{d}$ with $\mathbb{R}^{d}\times\{-1\}$, then according to [1], this projection also project unparametrized orbits of this spherical problem to unparametrized orbits of the two-center problem in $\mathbb{R}^{d}$, defined with respect to a modified metric $\|\cdot\|_{a}$ given below. The spherical energy thus induces an additional first integral for the two-center problem in $\mathbb{R}^{d}$ and vice versa, the energy of the Euclidean two-center problem is an additional first integral for the spherical two-center problem. Moreover, we may as well put a harmonic potential in the middle of the two Kepler centers and all these hold for the modified system, called the Lagrange problem [13] as well. This can as well be defined on $\mathbb{S}^{d}_{SH}$, with a potential proportional to the squared tangent of the central angle as the spherical harmonic potential.
We denote the masses of the two Kepler centers and the string constant of the Hooke center in $\mathbb{R}^{d}$ as $m_{1},m_{2},f\in\mathbb{R}$ respectively. We do not make assumptions on signs of them. We put the centers in $\mathbb{R}^{d}$ at $(0,\cdots,0,\pm a)$. Notice that in $\mathbb{R}^{d}$, the mechanical systems are defined with respect to the non-standard affine norm $\|(v_{1},\cdots,v_{d})\|_{a}:=\sqrt{v_{1}^{2}+\cdots+v_{d-1}^{2}+\frac{v_{d}^{2}}{1+a^{2}}}$.
@@
These systems, as well as those defined on $\mathbb{S}^{d}_{SH}$ are thus integrable. Similar result holds between systems defined in $\mathbb{S}^{d}$ and in $\mathbb{H}^{d}$ as well. We refer to [26] for precise statements. Note that setting $f=0$ in Prop 2.1 we obtain the relevant result in the two-center problem.
For the problem in $\mathbb{R}^{d}$, we denote the energy of the problem as $E_{Eucl}$. We consider $\mathbb{R}^{d}$ as a chart for the lower hemi-sphere $\mathbb{S}_{SH}$ and we denote by $E_{sph}$ the spherical energy in terms of positions and velocities, so as a function on the tangent bundle $T\mathbb{R}^{d}$. Note that Legendre transformation allows to identify the tangent and cotangent bundles, which then equips the tangent bundle with an induced symplectic form and an Poisson bracket.
@@ 2.3. Relationship among first integrals for the two center problem
We consider the Euclidean two center problem and argue in the case $d=2$. Due to the rotation invariance the formulas hold for any $d\geq 2$ as well. We follow the argument of [1] in this section. As mentioned in the introduction, we refer to [23] for an alternative report with full computational details.
We put the Kepler centers at $(0,a)$ and (0,a)(0,-a) in $\mathbb{R}^{2}$, with masses $m_{1}$ and $m_{2}$ respectively. We equip $\mathbb{R}^{2}$ with the norm $\|\cdot\|_{a}$, defined as $\|(u,v)\|_{a}=\sqrt{u^{2}+\frac{v^{2}}{1+a^{2}}}$.
This is the same setting as in [29], in which only one effective Kepler center is considered. By Prop 2.1 we may as well add the second Kepler center center to obtain the two-center problem. The energy of the Euclidean two-center problem is
@@
the energy of the corresponding (hemi-)spherical 2-center problem can be written, with the same kind of computation as in [29], as
@@
We use the affine coordinate change $(\xi,\eta)\mapsto(x=\xi,y=\sqrt{1+a^{2}}\eta+a)$ to bring the norm on $\mathbb{R}^{2}$ into the standard Euclidean one. For a quick comparison with [23] we remark that that such change transforms the unit sphere into an ellipsoid. With this change we have put the second center at $(0,-2a/\sqrt{1+a^{2}})$ in this new coordinates. And we have
@@
Further set $E_{Kep}:=E_{Eucl}|_{m_{2}=0}$ the Keplerian energy with respect to the first center. Then we can write
@@
in which
$D=(\xi\dot{\eta}-\eta\dot{\xi})^{2}-\dfrac{2a}{\sqrt{1+a^{2}}}\Bigl((\xi\dot{\eta}-\eta\dot{\xi})\dot{\xi}+\dfrac{m_{1}\eta}{\sqrt{\xi^{2}+\eta^{2}}}\Bigr)$
as appeared in (1) is a first integral of $E_{Kep}$: This is the first integral whose role we wish to address in this note. The term $V_{Eucl,2}$ is linear in $m_{2}$, as well as $K$, that we observe from its explicit expression
@@
The spherical energy $E_{sph}$ is conserved along orbits of the Euclidean two-center problem. Remind that we have used the Legendre transformation to identify the tangent bundle and the cotangent bundle. We may thus use the Legendre transformation to transport the Poisson bracket from the cotangent bundle to the tangent bundle for our convenience. This bracket is denoted still by $\{\cdot,\cdot\}$. We thus have
(5) $\{E_{Eucl},E_{sph}\}=0.$
@@ 2.4. Averaging in the two center problem
Remind that we have identified the tangent and cotangent bundles by means of Legendre transformation. We continue to work on the tangent bundle, which is equipped with a symplectic form and a Poisson-bracket induced from the cotangent bundle.
We set $m_{1}>0$ to ensure the existence of non-singular negative energy orbits of $E_{Kep}$. These orbits are all closed. We argue on the corresponding region of the tangent bundle, on which $E_{Kep}$ defines a Hamiltonian circle action and that these orbits in $\mathbb{R}^{d}$ avoid the second Kepler center. Denote the set of orbits of $E_{Kep}$ in $\mathbb{T}\mathbb{R}^{d}$ whose projection passes through the second Kepler center by $\mathcal{I}$. We shall exclude this set as $V_{Eucl,2}$ encounters singularities along these orbits. Then the region $\mathcal{R}$ is defined by having negative Keplerian energy $E_{Kep}<0$ and non-zero Kepler angular momentum $C_{Kep}\neq 0$ to exclude collisional orbits in $T\mathbb{R}^{d}\setminus\mathcal{I}$:
$\mathcal{R}:=\{E_{Kep}<0,C_{Kep}\neq 0\}\subset T\mathbb{R}^{d}\setminus\mathcal{I}.$
Remark 2.1.
The latter condition excluding collision orbits can likely be dropped since averaging is well-defined up to collisional orbits by mean of regularization. In dimension 2 and 3 this can be justified using methods and coordinates explained in [8] and [31] respectively. This is likely the case for all dimensions but we shall not deal with this issue here.
@@
$E_{Eucl,a}=E_{Kep}+\int_{S^{1}}V_{Eucl,2}\,ds,$
in which the integration is understood as averaging over closed orbits of $E_{Kep}$, parametrized by the circle $S^{1}=\mathbb{R}/\mathbb{Z}$.
The definition does not require the choice of a set of canonical coordinates. When $d=2,3$, several sets of canonical coordinates are available, including the classical Delaunay variables and Poincaré coordinates. Moser constructed action-angle coordinates of Delaunay type for $\mathbb{R}^{d}$ [18]. In general it is not always easy to construct these coordinates, which is even more complicated in the case when more bodies were involved [22]. Moreover, all these coordinates become singular at certain types of orbits.
Nevertheless, in a local trivialization of this circle bundle, we can have an (abstract) set of canonical coordinates $(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$ after a further localization on the space of closed Keplerian orbits, in which $L$ is the circular angular momentum as in the Delaunay coordinates, and $\tilde{\ell}$ is an angle parametrizing the Keplerian orbits in this trivialization (shifted from the mean anomaly by a phase, when both are well-defined), and $(u_{i},v_{i})$ are secular variables, in the sense that they are functions of the orbits only, independent of the actual position and velocity of the particle on the orbit. In [30] one finds detailed discussion on these type of abstract local coordinates.
Lemma 2.1.
Near any closed Keplerian orbit of $E_{Kep}$ in $\mathcal{R}$ there exists a set of canonical coordinates $(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$ such that $\tilde{\ell}$ is an angle parametrizing the closed Keplerian orbits, $L$ is its conjugate action variable on which $E_{Kep}$ only depends, and $(u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$ are secular coordinates which depends only on the Keplerian orbits.
Proof.
In a local trivialization we have the angle $\tilde{\ell}$ and its conjugate action $L$ (as a moment map of the Hamiltonian circle action). If we symplectically reduce a local trivialization of this circle bundle with respect to the Hamiltonian circle action, then by Darboux theorem there is a set of symplectic coordinates $(\tilde{u}_{1},\cdots,\tilde{u}_{d-1},\tilde{v}_{1},\cdots\tilde{v}_{d-1})$. Then we recover that $(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$ is a set of canonical coordinates, in which $u_{i}=f_{i}(L)\tilde{u}_{i},v_{i}=g_{i}(L)\tilde{v}_{i}$ for some non-zero functions $f_{i},g_{i}$ of $L$. Indeed these functions cannot depend on $\tilde{\ell}$ as a consequence of the unicity of the moment map of a Hamiltonian circle action up to a constant. It is possible to say more about the functions $f_{i},g_{i}$ but we shall not need to do this for our current purpose. ∎
It is not hard to see that Definition 2.1 is independent of the choice of local coordinates. Thus averaging is well-defined in a non-trivial principal $S^{1}$-bundle, in our case defined by motions along closed Keplerian orbits.
Lemma 2.2.
The expression $\{G_{1}(L),G_{2}(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})\}$ has zero-average over a period of $\tilde{\ell}$.
Proof.
In coordinates,
$\{G_{1}(L),G_{2}(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})\}=G_{1}^{\prime}(L)\,\dfrac{d}{d\tilde{\ell}}(G_{2}).$
Integration of a total derivative over a period of $\tilde{\ell}$ is zero, and thus the conclusion. ∎
The following proposition is a coordinate-free formulation of a main observation in [19]. Without specifying the canonical coordinates used, the dimension of the space and the sign of $m_{2}$ in the announcement we bring a small generalization.
@@
We cover $M$ by open submanifolds on which the above-mentioned canonical coordinates are defined. Let $\tilde{M}$ be any of these submanifolds, with local canonical coordinates $(L,\tilde{\ell},u_{1},\cdots,u_{d-1},v_{1},\cdots v_{d-1})$ as described in Lem 2.1.
@@
$E_{Eucl,a}=E_{Kep}+\langle V_{Eucl,2}\rangle_{\tilde{\ell}}:=E_{Kep}+\dfrac{1}{2\pi}\int_{0}^{2\pi}V_{Eucl,2}d\tilde{\ell}$.
Then
$E_{Eucl}=E_{Eucl,a}+V_{Eucl,2}-\langle V_{Eucl,2}\rangle_{\tilde{\ell}}$.
We apply averaging theory. There is only one frequency that we need to average. Moreover we may use $m_{2}$ to play the role of a small parameter. In this setting, we know that with one step of averaging we get a canonical transformation $\phi:\tilde{M}\mapsto\tilde{M}$ of order $O(m_{2})$ which conjugate $E_{Eucl}$ to $E_{Eucl,a}$ up to an error term of order $O(m_{2}^{2})$, that is to say,
$E_{Eucl,a}\dot{=}E_{Eucl}\circ\phi.$
@@
$\{E_{Eucl,a},D\circ\phi+K\circ\phi\}\dot{=}0.$
@@
$\{E_{Eucl,a},D\circ\phi+K\circ\phi\}\dot{=}0=\{E_{Eucl,a},D\}+O(m_{2}).$
The expression $\{E_{Eucl,a},D\}$ is independent of $\tilde{\ell}_{1}$, as well as its term of order $m_{2}$. The contribution of order $m_{2}$ in the remainder term $O(m_{2})$ comes from $\{E_{Kep},D\circ\phi+K\circ\phi\}$, which has zero-average while integrated with respect to $\tilde{\ell}_{1}$ over $(0,2\pi)$ according to Lem 2.2. So by integration we obtain
$\{E_{Eucl,a},D\}\dot{=}0.$
and thus
$\{E_{Eucl,a},D\}=0,$